### v3

# VIPTR: A Vision Permutable Extractor for Fast and Efficient Scene Text Recognition

Abstract
Scene Text Recognition (STR) is a challenging task that involves recognizing text within images of natural scenes. Although current state-of-the-art models for STR exhibit high performance, they typically suffer from low inference efficiency due to their reliance on hybrid architectures comprised of visual encoders and sequence decoders. In this work, we propose the VIsion Permutable extractor for fast and efficient scene Text Recognition (VIPTR), which achieves an impressive balance between high-performance and rapid inference speeds in the domain of STR. Specifically, VIPTR leverages a visual-semantic extractor with a pyramid structure, characterized by multiple self-attention layers, while eschewing the traditional sequence decoder. This design choice results in a lightweight and efficient model capable of handling inputs of varying sizes. Extensive experimental results on various standard datasets for both Chinese and English scene text recognition validate the superiority of VIPTR. Notably, the VIPTR-T (Tiny) variant delivers highly competitive accuracy on par with other lightweight models and achieves SOTA inference speeds. Meanwhile, the VIPTR-L (Large) variant attains greater recognition accuracy, while maintaining a low parameter count and favorable inference speed. Our proposed method provides a compelling solution for the STR challenge, which blends high accuracy with efficiency and greatly benefits real-world applications requiring fast and reliable text recognition. The code is publicly available at https://github.com/cxfyxl/VIPTR.
1 Introduction
Text Recognition (TR) is crucial for multimodal document analysis, autonomous driving, office automation, and other applications Shi et al. (2016); Guo et al. (2022); Yang et al. (2020); Guo et al. (2023); Li et al. (2023a); Wu et al. (2023); Zhou et al. (2023). Recent deep learning advancements have spurred growth in Scene Text Recognition (STR), which aims to digitize text from images in natural settings for use in language processing tasks. Despite progress, STR still faces hurdles such as distorted text, varied fonts, shadows, and complex backgrounds. Efforts are ongoing to improve recognition precision, with an emphasis on creating models that are accurate, fast, and lightweight for versatile deployment.
Figure 1: Model Architecture Evolution for Scene Text Recognition.
Scene text recognition initially relied on deep learning models like CRNNs Shi et al. (2016) that combined visual feature extraction with sequence processing, favored for their fast inference despite sensitivity to text disturbances. Subsequent models adopted encoder-decoder, autoregressive approaches, enhancing accuracy through contextual understanding but at the cost of slower word-by-word translation speeds. More recent advances involve visual semantic parsing frameworks that integrate semantic knowledge to improve recognition accuracy in specific contexts, which often require large models or intricate semantic systems, leading to efficiency compromise. Recent researchers 2021 (2020); Zhang et al. (2023) tried to develop simple and effective model structures by using the attention mechanism to aggregate the sequence features corresponding to the same character, in addition, align the CE loss and CTC loss to improve the convergence speed of the CTC loss. From the perspective of simplifying the model architecture, some models used only CNN Borisyuk et al. (2018) or ViT Atienza (2021) for identification, which is efficient in inference but has a large gap in accuracy compared to SOTA methods. Inspired by the pyramid structure of Swin Transformer Liu et al. (2021), the SVTR model Du et al. (2022), designed for single visual modality input, is proposed to simultaneously process the implicit local visual information in text images and unsupervised mining of global contextual semantic information. SVTR has four structural variants with different capacities.
The STR models based on vision Transformer set a typical example for fast and effective STR schemes for a single visual modality. On one hand, it practices the effectiveness of replacing convolution with a self-attention mechanism to realize a new visual feature extractor. On the other hand, it reveals the importance of modeling global dependencies of lateral symbol sequences in text image recognition tasks, which even replace bidirectional LSTMs. Despite the outstanding performance of current ViT-based STR, it is clear that it still has several shortcomings: (1) The computation of self-attention relies on absolute position encoding and fixed-size mask operation, which limits the length of the input image; (2) In each stage, only one type of global and local relations is mined separately, and the distance between the steps of global dependencies is limited to save reasoning time, so the recognition effect of long text images is not good; (3) The actual inference time consumption of the model in general scenarios is still larger than that of CRNN with the same parameter magnitude.
To further enhance the vision model and meticulously control computational complexity to achieve more efficient STR while maintaining high accuracy, we propose VIPTR—a vision-based approach for a faster, more accurate, and cross-language universal STR model. Inspired by Cross-Shaped Windows Vision Transformer (CSwin) Dong et al. (2022) and other visual models based on enhanced self-attention mechanisms, VIPTR first divides the text image evenly into multiple 2D patches, each patch contains part of the characters or background. This way of dividing patches facilitates the simultaneous extraction of local features of characters and the capture of potential contextual semantic dependencies between different characters. In each token mixing attention module, we try to use conditional positional embedding (CPE) Chu et al. (2021) combined with Manhattan self-attention (MaSA) Fan et al. (2023) and CSWin Attention to explore the local information of characters, and use Overlapping Spatial Reduction Attention (OSRA) Lou et al. (2023) module to model global dependencies. Specifically, along the horizontal direction Model the global dependency between characters in the axial direction, model the global relationship from top to bottom of the same character along the vertical axis, and align to achieve weight sharing. Compared with SWin, this method significantly reduces the time complexity and does not require the use of a fixed-size mask group to limit the length of the input image. At the same time, we also use a pyramid structure for multi-stage local and global information hybrid modeling, which greatly improves the feature extraction capabilities of text images. To sum up, VIPTR can surpass the recognition effect and inference efficiency of SVTR at a smaller model size.
The main contributions of this work can be summarized as.
  • We verified that the single-vision model based on the self-attention mechanism can still achieve comparable accuracy to the high-level vision-language model in the scene text recognition task. At the same time, the use of sparse operators and different attention combinations can accelerate the calculation of the self-attention mechanism, and achieve the balance between performance and speed advantages that STR has been pursuing.
  • We proposed VIPTR, a feature extraction module tailored for parsing text. The module can adapt to and accurately recognize the text image input of any length and has the versatility of cross-language recognition, which has a good application prospect.
  • We verified the superiority of VIPTR on cross-language benchmark datasets and manually annotated industrial application datasets. Among them, VIPTRv1-L achieved SOTA in recognizing Chinese and English scene text. VIPTRv2-T achieves the most efficient inference while ensuring accuracy. It is inferred on an NVIDIA V100 GPU with a parameter of 5.1M and consumes an average of 3.3ms per text image.
2 Related Work
Figure 2: Overall architecture of the proposed VIPTR. It is a four-stage network with three-stage height progressively decreasing. In each stage, a series of attention-mixing blocks are carried out and followed by a subsampling or combining operation. In particular, VIPTR designs two visual feature fusion modes in vision permutation blocks: series and parallel. At last, the recognition is conducted by the CTC decoder.
CNN-Based Vision Modules
Convolutional neural networks (CNN) serve as pivotal tools for extracting visual information in previous tasks. VGG Simonyan and Zisserman (2014) demonstrated that a series of 3x3 convolutions is sufficient to train a high-performance model. Later, ResNet He et al. (2016) introduced residual connections along with the batch normalization (BN) layer, which enabled the training of exceedingly deep neural networks and further improved performance. In the realm of STR, the CNN-based vision model transforms a pyramid structure capable of processing local features across multiple stages. This transformation effectively compresses the two-dimensional image input into a one-dimensional feature sequence, allowing for decoding the sequence without reliance on recurrent neural networks (RNN). High recognition accuracy can also be obtained, which proves the effectiveness of locality bias extracted by CNN. We believe that the experience gained in CNN-based vision models, including the pyramid structure for extracting locality bias through small kernel convolutions, residual connections, and step-by-step downsampling, remains valid in current research. Therefore, these techniques continue to be utilized in our proposed STR architecture.
Transformer-Based Vision Modules
ViT Dosovitskiy et al. (2020) is the first work to introduce Transformer Vaswani et al. (2017); Yang et al. (2022); Touvron et al. (2021) into the construction backbone of vision tasks. Experiments demonstrate that, with sufficient training data, ViT outperforms traditional CNNs in vision tasks. DeiT Touvron et al. (2021) introduces a new training strategy, enabling ViT to achieve superior performance on small-scale training sets such as ImageNet-1K. These methods demonstrate the efficacy of global dependency modeling through self-attention mechanisms. Later, the pyramid structure also enhances the performance of Transformer-based models across various vision tasks. Based on the pyramid structure, Swin Transformer proposes to use self-attention within each local window. This can be regarded as a use of locality bias. Vision Transformers are integrated into STR tasks as visual feature encoders, leading to a significant improvement in accuracy. ViTSTR Atienza (2021), visionLAN Wang et al. (2021b), and MaskOCR Lyu et al. (2022) have all come close to the SOTA effect. SVTR-T both ensures high performance and creates the fastest recognition efficiency. Although the self-attention mechanism has greatly improved the effect of text recognition, its computational complexity is quadratic compared to the number of input tokens. As the resolution of the input image increases, the inference time consumption also increases quadratically.
Efficient Self-Attention Based Vision Modules
A challenging problem in the design of Transformer is that the computational cost of global self-attention is very expensive Vaswani et al. (2017); Yang et al. (2021, 2023), while local self-attention usually limits the interaction field of each token, resulting in insufficient information mining. The CSWin Self-Attention mechanism uses parallel computing to form the self-attention of the horizontal and vertical stripes of the cross-shaped window, achieving powerful feature modeling capabilities while limiting computing costs. Decomposed MaSA achieves efficient and significant feature representation by defining a spatial attention attenuation matrix and its decomposition paradigm based on two-dimensional Manhattan distance, effectively layering local information on a global scale. Spatial reduction attention (SRA) Wang et al. (2021a) effectively extracts global information by utilizing sparse token-region dependencies, and OSRA introduces overlapping spatial reduction (OSR) Lou et al. (2023) into SRA to achieve better performance by using larger and overlapping patches.
3 Method
3.1 Design Guidelines
In this study, we use some important properties of CNNs, ResNet, and several self-attention mechanisms in a new architecture inspired by hierarchical vision Transformers to design a paradigm of high-performance networks for STR. Here are the design guidelines we followed:
  1. 1.
    Explicitly inject character features into the front of the network in the form of locality bias.
  2. 2.
    Design a new permutation method to effectively fuse the character component features of the Local and Global Attention modules in the middle layer.
  3. 3.
    Implicitly handle image aspect and width resolution in the network initialization structure, canceling or replacing all modules that depend on the input size.
3.2 Overall Architecture
The schematic representation of the proposed VIPTR is depicted in Figure 2. It is a four-stage network with three height progressively decreased layers. Given an image text of size $H\times W\times 3$, it undergoes an initial transformation into $\frac{H}{4}\times\frac{W}{4}\times C_{0}$ patches through a two-layer CNN-based patch embedding. Since each patch is associated with a text character or part of the background in the image, these patches are called character components. Subsequently, four stages of feature extraction are performed at different scales, each stage consists of a series of self-attention-based hybrid blocks, followed by downsampling or global average pooling operations along the height dimension. Both Local and Global Attention blocks are employed to concentrate on character feature extraction and capture dependencies between character components. The extractor serves to characterize character component features at different distances and multi-scales and the contextual dependencies between character components, and a representation called $B$ with a size of $1\times\frac{W}{4}\times C_{3}$ is generated, which captures multi-granularity character features. Finally, the CTC decoder is used to perform linear prediction and deduplication on the character sequence to obtain the recognition result.
3.3 Patch Embedding
The first step of the vision Transformer is to cut the image into several equal-sized patches, and then hand them over to the subsequent self-attention module for processing in the form of several sets of sequences. In this article, for an image text, first convert it from $X\in R^{H\times W\times 3}$ into several character components $B_{0}B_{1}\in R^{\frac{H}{4}\times\frac{W}{4}\times C_{0}}$. Alternatively, we implement the patch embedding by using two consecutive $3\times 3$ convolutions with stride $2$ and BNs. This scheme Du et al. (2022) gradually increases feature dimensions, which is beneficial to feature fusion.
Figure 3: Illustration of (a) Cross-Shaped Windows Self-Attention (CSWin), (b) Decomposed Manhattan Self-Attention (D-MaSA), (c) Multi-Head Self-Attention (MHSA), and (d) Overlapping Spatial Reduction Attention (OSRA).
3.4 Character and String Modeling
Existing studies mainly use feature sequences with height $1$ to represent image text. Each feature corresponds to a small local region. However, these segments are usually noisy, especially for irregular text, failing to describe the characters of an image well. In contrast, the Vision Transformer introduces a two-dimensional feature representation, which enables more accurate discrimination between different characters and backgrounds. This suggests that two-dimensional features can be used to effectively represent the local strokes of characters and the potential contextual dependencies of strings.
More specifically, image text requires features to be extracted from two perspectives. The first involves local character features, including elements like strokes and fonts, representing internal associations of a character and morphological features. The second is context relations, such as correlations between different characters or between text and non-text components. Therefore, we design two classes of attention-mixing blocks with different receptive fields to sense relevance. Mathematically speaking, for a sequence of character components $B_{i-1}B_{i}$ in the previous stage, it is first reshaped as a sequence of features. Upon the input of the attention mixing block, Layer Normalization (LN) and Position Embedding (PE) are first applied. Subsequently, the self-attention mechanism is employed for dependency modeling. Finally, LN and Feed Forward Network (FFN) are successively applied for feature fusion, with the residual connection applied.
Positional Embedding.
Since self-attention is permutation invariant and ignores tokens’ positional information, positional Embedding (PE) is widely used in Transformers to add such information back. Typical positional Embedding mechanisms include absolute positional encoding (APE) Vaswani et al. (2017), relative positional encoding (RPE) Liu et al. (2021), and CPE. APE and RPE are designed for a specific input size and are not friendly to varying input resolutions. CPE takes the feature as input and can generate the positional encoding for arbitrary input resolutions. Then the generated positional encoding will be added to the input feature. Locally-enhanced Positional Encoding (LePE) Dong et al. (2022) shares a similar spirit as CPE, but proposes to add the positional encoding as a parallel module to the self-attention operation and operates on projected values in each Transformer block.
Local Attention.
The goal of Local Attention is to encode character morphological features and establish correlations within character components, thereby simulating stroke features. Local Attention needs to consider the neighborhood of character components on the two-dimensional coordinate system. Drawing on the design ideas of MaSA and CSWin Self-Attention, we designed two efficient self-attention mechanisms that explicitly model local features on a two-dimensional coordinate system. MaSA is shown in Figure 3LABEL:sub@subfig:dmsa. The spatial attenuation matrix of the image means that character features are highlighted globally and background features are weakened. CSWin Self-Attention is shown in Figure 3LABEL:sub@subfig:cswin. This mechanism is implemented by executing local self-attention of horizontal and vertical stripes of step size called Split-Window ($SW$) decomposed from a certain cross window in parallel. For image text input, considering that the image resolution is larger than the Channel in the early stage, we choose a smaller $SW$ in the early stage, and because the size of the input image text changes dynamically during processing, and the length and width resolution needs to be guaranteed. Divisible by $SW$, we empirically set $SW$ for stages using CSWin Self-Attention to $1$ or $2$ by default.
Global Attention.
As shown in Figure 3LABEL:sub@subfig:mhsa, Global Attention calculates the contextual dependencies between all character components. Since text and non-text are the two main elements in an image, this universal blend can create long-distance dependencies between the components of different characters. In addition, using the self-attention mechanism can also weaken the influence of non-text components while enhancing the importance of text components. In addition, to save computational costs, we introduced OSRA, as Figure 3LABEL:sub@subfig:osra, which adds overlapping convolution operations to MHSA to sparse character components in some channels into more salient features without losing edge information. The use of salient features greatly reduces the number of tokens required to calculate the self-attention mechanism, thus improving computational efficiency and performance.
As mentioned above, the hybrid blocks in VIPTR are designed to extract distinct yet complementary features. In this way, blocks are initially processed locally before being addressed globally. At each stage, Local and Global Attention are recurrently applied multiple times for comprehensive feature extraction. Effect of the two kinds of blocks Permutations can be seen in ablation experiments.
3.5 Subsampling
To save computational costs and extract significant features layer by layer, we added a down-sampling operation Du et al. (2022) on the height dimension after the first and third attention mixing stages of VIPTR. Using the output of the features from the previous Attention block, we first reshape it into an embedding of size $H\times W\times C_{i-1}$, representing the current height, width, and channel respectively. We then employ a $3\times 3$ convolution with a height dimension stride $2$ and a width dimension stride $1$, followed by LN, producing an embedding of size $\frac{H}{2}\times W\times C_{i}$. The Subsampling block halves the height while maintaining a constant width. It not only reduces the computational cost, but also compresses the height dimension to establish a multi-scale representation for each character without affecting the token layout in the width dimension. We also increase the channel dimension $C_{i}$ to compensate for the information loss.
3.6 Combining and Prediction
In the final stage, we use combinatorial operations Du et al. (2022) to integrate sequence features. It first uses AdaptiveAvgPool to compress the height dimension to $1$, followed by fully connected layers, Hardswish activation, and dropout. By doing so, the character components are further compressed into feature sequences, where each element is represented by a feature of length $C_{3}$. Using AdaptiveAvgPool avoids applying convolution to very small embeddings of one dimension, such as height $2$, while being able to apply a variety of initial input heights.
We finally use the CTC decoder to perform parallel linear prediction on the obtained combined features to achieve recognition. Specifically, a linear classifier with $N$ nodes is adopted. It generates transcribed sequences of size $W/4$, where components of identical characters are ideally transcribed as repeating characters and components of non-text are transcribed as whitespace symbols. The sequence is automatically compressed into the final result. For example, $N$ is set to $37$ for English when implemented.
3.7 Vision Permutation and Architecture Variants
Firstly, as shown in Figure 2, in the 2-3 stages of VIPTR, we design the structures of Local and Global Attention to process character components circularly in two ways: ”series processing” and ”block-parallel processing-merging”, respectively, called VIPTRv1 and VIPTRv2. Second, we design several hyperparameters in VIPTR, including channel depth and number of heads per stage, and number of mixing blocks. By varying them, VIPTR architectures with different numbers of parameters can be obtained. In summary, we constructed four typical architectures, namely VIPTRv1-T (Tiny), VIPTRv2-T, VIPTRv1-L (Large), and VIPTRv2-B (Base). Its detailed configuration is shown in Table 2.
Atten. Permutation IC13 IC15 FLOPs (G)
CRNN (None) 75.3 56.7 0.69
SVTR-T 77.9 62.5 0.29
$[L_{1}][L_{1}G_{1}][G_{1}]$ 79.1 65 0.269
$[L_{1}][L_{1}G_{2}][G_{1}]$ 79.1 65.2 0.258
$[L_{2}][L_{2}G_{1}][G_{1}]$ 79 65.1 0.282
$[L_{2}][L_{2}G_{2}][G_{1}]$ 79.3 65.3 0.27
$[L_{1}][L_{2}G_{1}][G_{1}]$ 78.8 64.9 0.27
$[L_{2}][L_{1}G_{1}][G_{1}]$ 78.9 65.1 0.27
$[L_{1}][L_{2}G_{2}][G_{1}]$ 78.7 65 0.258
$[L_{2}][L_{1}G_{2}][G_{1}]$ 78.8 64.8 0.258
$[L_{1}][L_{1}//G_{1}][G_{1}]$ 75.5 60.5 0.195
$[L_{1}][L_{1}//G_{2}][G_{1}]$ 75.4 60.5 0.192
$[L_{2}][L_{2}//G_{1}][G_{1}]$ 75.6 60.4 0.197
$[L_{2}][L_{2}//G_{2}][G_{1}]$ 75.8 61 0.194
$[L_{1}][L_{2}//G_{1}][G_{1}]$ 75 59.9 0.196
$[L_{2}][L_{1}//G_{1}][G_{1}]$ 75.2 59.7 0.196
$[L_{1}][L_{2}//G_{2}][G_{1}]$ 74.9 60.1 0.193
$[L_{2}][L_{1}//G_{2}][G_{1}]$ 75.2 60.3 0.193
Table 1: Ablation study on Attention Blocks’ permutations.
Model $[C_{0},C_{1},C_{2},C_{3}]$ $[N_{1},N_{2},N_{3},N_{4}]$ Heads Attention Permutation Params (M) FLOPs (G)
VIPTRv1-T $[64,128,256,192]$ $[3,3,3,3]$ $[2,4,4,8]$ $[L_{1}][L_{1}G_{2}][G_{1}]$ 4.0 0.26
VIPTRv1-L $[192,256,512,384]$ $[3,7,2,9]$ $[6,8,8,16]$ $[L_{2}][L_{2}G_{2}][G_{1}]$ 37.7 2.31
VIPTRv2-T $[64,128,256,192]$ $[3,3,3,3]$ $[2,4,4,8]$ $[L_{1}][L_{1}//G_{2}][G_{1}]$ 3.2 0.19
VIPTRv2-B $[128,256,384,256]$ $[3,6,6,9]$ $[4,8,8,12]$ $[L_{2}][L_{2}//G_{2}][G_{1}]$ 20.2 1.18
Table 2: VIPTR architecture variants (w/o counting the rectification module and CTC decoder).
4 Experiments and Discussion
Datasets
Our English STR models are trained on two synthetic datasets, MJSynth (MJ) Jaderberg et al. (2014, 2016) and SynthText (ST) Gupta et al. (2016) trainsets (72k+55k), along with real-world data from COCO-Text Gomez et al. (2017). Testing is conducted on six public benchmarks: ICDAR 2013 (IC13) Karatzas et al. (2013) testset with 857 regular images, Street View Text (SVT) Wang et al. (2011) testset with 647 images from Google Street View, IIIT5K-3000 Mishra et al. (2012) sourced from websites, ICDAR 2015 (IC15) Karatzas et al. (2015) comprising 1811 images captured via Google Glasses, Street View Text Perspective (SVTP) Phan et al. (2013) with 639 images also from Google Street View, and CUTE80 (CUTE), which focuses on curved text recognition, with 288 images.
For the Chinese recognition task, we employ the Chinese Scene and Document Dataset Yu et al. (2021). The former includes a scene text dataset with 509,164 training, 63,645 validation, and 63,646 test images, featuring 5,880 unique characters. The document text dataset comprises 400,000 training, 50,000 validation, and 50,000 test images, with 4,865 distinct characters.
Implementations
The recommended training and testing configurations for VIPTR are listed in Table 3. When Training English models, data augmentation like stretch, perspective distortion, blur, and Gaussian noise, are randomly performed with probability $0.4$. Then, It was not used when training the Chinese model. All models are trained by PyTorch on four Tesla V100 GPUs.
Setup English model Chinese model
Rectification module Shi et al. (2018)
Distortion correction $32\times 64$
Optimizer AdamW
Weight decay 0.05
Lr scheduler Cosine annealing
Input size $32\times 96$ $32\times 320$
Batch size 1024 512
Initial learning rate 5e-4×$\times\frac{batchsize}{2048}$ 3e-4×$\times\frac{batchsize}{512}$
Training epochs 49 100
Warm-up epochs 2 5
Evaluation metric Word accuracy
Table 3: Training configurations for VIPTR.
Comparison with State-of-the-Art
As indicated in Table 6, we compare VIPTR with prior studies on six English and one Chinese scene benchmarks. Even the fastest variant, VIPTRv2-T, achieves highly competitive accuracy, particularly among models without Language Models (LM). In English datasets, VIPTRv1-L outperforms LM-used models and SVTR-L, even achieving the best accuracy on IC15. Both VIPTRv1-L and VIPTRv2-B demonstrate overall accuracy comparable to recent studies Fang et al. (2021); Tang et al. (2021); Du et al. (2022); Lyu et al. (2022) while maintaining simplicity and faster execution.
Turning to the Chinese Scene Dataset, the accuracy of 8 existing methods is provided by Du et al. (2022). Encouragingly, VIPTR performs notably well, with a 0.5% accuracy gain compared to SVTR-L, the best-performing model among those listed. Other VIPTR variants also exhibit advanced Chinese recognition accuracy. These results are attributed to VIPTR’s superior ability to perceive multi-grained character component features and its breakthroughs in characterizing Chinese words with rich stroke patterns.
4.1 Further Analysis
Ablation Experiments
To gain a deeper insight into VIPTR, we conducted controlled experiments on both IC13 (regular) and IC15 (irregular) under various configurations. The English dataset used in the experiment consists of COCO-Text and part of the real scene training set, with a total of 45054 images over 500 epochs of training. In addition, we conduct a series of comparative experiments to verify the effectiveness of our models on image texts of varying lengths. The Chinese Document Dataset is used for training and validation. Three test sets are employed: Testset1, the standard set from the Chinese Document Dataset; Testset2, consisting of 2,367 images cropped from real documents with lengths between 320 and 1500 pixels; and Testset3, featuring 638 images, also derived from real documents but with lengths exceeding 1500 pixels. For efficiency, all the experiments are carried out by using VIPTR-T without the rectification module and data augmentation.
Positional Embedding IC13 IC15 FLOPs (G)
None 76.7 59.7 0.25
APE 77.3 63.5 0.26
CPE 77.3 63.6 0.26
LePE 79.1 65 0.26
Table 4: Ablation study on Positional Embeddings.
Model Testset1 Testset2 Testset3
CRNN 97.5 91 91.6
SVTR-T 98.2 47.5 5.9
VIPTRv1-T 98.8 92.1 93.7
Table 5: Ablation studies on inputs of different lengths.
Methods English Regular English Unregular Avg. Chinese Scene Params (M) Speed $\downarrow$ (ms)
IC13 SVT IIIT5k IC15 SVTP CUTE
CRNN Shi et al. (2016) 91.1 81.6 82.9 69.4 70.0 65.5 78.5 53.4 8.3 6.3
Rosetta Borisyuk et al. (2018) 90.9 84.7 84.3 71.2 73.8 69.2 80.3 - 44.3 10.5
ViTSTR Atienza (2021) 93.2 87.7 88.4 78.5 81.8 81.3 85.6 - 85.5 11.2
SVTR-T Du et al. (2022) 96.3 91.6 94.4 84.1 85.4 88.2 90.7 67.9 6.0 4.5
SVTR-L Du et al. (2022) 97.2 91.7 96.3 86.6 88.4 95.1 92.8 72.1 40.8 18.0
ASTER Shi et al. (2018) 90.8 89.5 93.4 76.1 78.5 79.5 86.5 54.5 27.2 -
NRTR Sheng et al. (2019) 94.7 88.3 86.5 74.1 79.4 88.2 84.0 - 31.7 160.0
SAR Li et al. (2019) 91.0 84.5 91.5 69.2 76.4 83.5 83.6 62.5 57.5 120.0
SATRN Lee et al. (2020) 95.7 93.5 96.1 84.1 88.5 90.3 91.9 - - -
SRN Yu et al. (2020) 95.5 91.5 94.8 82.7 85.1 87.8 90.4 60.1 54.7 25.4
ABINet Fang et al. (2021) 97.4 93.5 96.2 86.0 89.3 89.2 92.6 - 36.7 51.3
VisionLAN Wang et al. (2021b) 95.7 91.7 95.8 83.7 86.0 88.5 91.2 - 32.8 28.0
VST Tang et al. (2021) 96.4 93.8 96.3 85.4 88.7 95.1 92.6 - 64.0 -
MATRN Na et al. (2022) 95.8 94.9 96.7 82.9 90.5 94.1 92.3 - 44.2 29.6
MaskOCR* Lyu et al. (2022) 97.7 93.7 95.4 86.6 89.0 87.5 92.5 68.6 36.0 24.1
MaskOCR Lyu et al. (2022) 96.8 94.7 95.3 87.1 89.3 90.6 92.7 68.8 100.0 -
TrOCR Li et al. (2023b) 98.3 93.2 91.0 84.0 91.0 89.6 90.2 - 558.0 20.8
AutoSTR Zhang et al. (2020) 94.2 90.9 94.7 81.8 81.7 84.0 89.5 - 10.4 207.0
PREN* Yan et al. (2021) 94.7 92.0 92.1 79.2 83.9 81.3 88.0 - 29.1 40.0
VIPTRv1-T (Tiny) 96.8 93.5 93.3 87.2 87.1 87.4 91.4 68.7 5.9 4.2
VIPTRv2-T 95.8 92.6 92.6 85.4 86.0 85.0 90.3 65.8 5.1 3.3
VIPTRv1-L (Large) 97.3 94.5 95.1 88.4 89.6 88.8 92.9 72.6 39.5 16.9
VIPTRv2-B (Base) 97.0 94.2 94.8 88.0 90.0 90.2 92.7 71.0 22.1 8.0
Table 6: Results on six English and Chinese scene benchmarks tested against existing methods. The model types from top to bottom are CTC-based, attention-based, LM-based, and others. * means the lightweight version of the corresponding model. The speed is the average inference time on one Tesla V100 GPU with batch size 4 (to align with Du et al. (2022)) for 3000 English image texts.
Effective Permutation of Attention Blocks
There are various ways to match the global and Local Attention blocks in each stage. In previous researchv Du et al. (2022), it was found that by placing the Local Attention block before the Global Attention block, it is beneficial to both feature mining of individual symbols and to guide the Global Attention block to focus on long-term dependency capture between symbol features, while effectively saving the inference cost of the model. The top half and the bottom half of Table 1 respectively show the performance of VIPTRv1-T and VIPTRv2-T composed of various Local and Global Attention and their permutations. $[L_{i}][L_{j}G_{s}][G_{t}]$ means for each stage, Local Attention blocks are carried out at first, and then Global Attention blocks. $[L_{i}][L_{j}//G_{s}][G_{t}]$ means in the stage 2-3, Local and Global Attention blocks are carried in parallel. The $G_{1}$ means MHSA, the $G_{2}$ means ORSA, the $L_{1}$ means CSWin and the $L_{2}$ means MaSA. Others are defined similarly.
It can be observed that almost every scheme gives a certain accuracy improvement compared to CRNN. We believe that the improvements are attributed to the comprehensive perception of character components features by the self-attention mechanism. The relatively large gains on irregular text further explain the self-attention mixing block is helpful to feature learning in complex scenarios. It is observed that $[L_{2}][L_{2}G_{2}][G_{1}]$ reports the best accuracy in the top half of Table 1 and $[L_{2}][L_{2}//G_{2}][G_{1}]$ achieves best in the bottom half of Table 1. The best one gives accuracy gains of 1.4% and 2.8% on IC13 and IC15 when compared with SVTR-T. It is observed that $[L_{1}][L_{1}//G_{2}][G_{1}]$ reports the fastest inferencing speed.
4.1.1 Effectiveness of Positional Embeddings.
As indicated in Table 4, various PE schemes exhibit slight variations in recognition accuracy. The LePE scheme surpasses the default one (None) by 2.4% and 5.3% on the two datasets, demonstrating its effectiveness, particularly in the case of irregular text.
4.1.2 Results for different lengths inputs testing
In Table 5, the performance of the three models on the three test sets is quite different. SVTR’s inference performance starts to collapse as the input length gets larger. VIPTR not only supports arbitrary lengths of image text input, but the inference performance will not decrease as the input length increases.
5 Conclusion
In this study, we propose VIPTR, an innovative vision permutable extractor for fast and efficient Scene Text Recognition (STR). VIPTR excels at extracting character features and capturing contextual dependencies between character components, all while being unrestricted by input size during inference. These grants VIPTR notable accuracy, efficiency, and cross-lingual adaptability. We have also developed different versions of VIPTR to cater to a range of application requirements. Experiments on both English and Chinese benchmarks generally validate the effectiveness of our models. It achieves highly competitive or superior accuracy compared to state-of-the-art methods while maintaining faster processing speeds. Our future goal is to integrate VIPTR with an efficient text decoder to further elevate STR performance.